When a parachute opens, the air exerts a large drag force on it.This upward force is initially greater than the weight of the skydiver and, thus, slows him down. Suppose the weight of the skydiver is 915 N and the drag force has a magnitude of 1027 N. Themass of the sky diver is 93.4 kg. What are the magnitude anddirection of his acceleration?
Answer #1
total force acting on skydiver?    1027 up plus 915 down gives...
    1027 - 915 = 112 N  up
And using newtons second law:    F =ma     or     a =F/m = 112 / 93.4 = 1.20 m/s2 upward